OpenStudy (anonymous):

Describe how the graph of y = x2 can be transformed to the graph of the given equation. y = (x - 2)2 - 15

OpenStudy (debbieg):

Refer to the answer I gave you here: http://openstudy.com/study#/updates/5223c882e4b06211a67f5c85 And also: y=f(x)+k will vertically translate the graph of y=f(x) UP k units if k>0, or DOWN |k| units if k<0. E.g., y=x^2 - 2 is a vertical shift down of y=x^2. Now combine that with the horizontal translation stuff from the earlier question, and you should be able to answer this one.

OpenStudy (anonymous):

Shift the graph of y = x2 left 2 units and then down 15 units.

OpenStudy (debbieg):

Down 15 is correct. Be careful with the direction of the horizontal shift... go back to the other thread to review the way that h affects the direction. Or notice: what x is an x-intercept for \(y=x^2\) ? What about for \(y=(x-2)^2\) ? Which direction does the vertex move?